the sum of two numbers is 208. one-third of the greater numbers equals the smaller number. what are the numbers?
Answer by mananth(16946)

the sum of two numbers is 208. one-third of the greater numbers equals the smaller number. what are the numbers?
Let one number be x
The other number will be 208-x
1/3 *x = 208-x
1/3 x + x = 208
4/3 *x = 208
4x=208*3= 624
x=156
208-156 = 52
The numbers are 156 and 52